News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 16 July 2004 11:30am EDT

During the last week Chandra completed the observing schedule as planned.

A real-time procedure was completed on July 15 to the contents of ACIS Back-end Processor (BEP) cache, EEPROM, and bulk memory, warm-boot the BEP, and re-dump. This procedure was to gather data to aid in diagnosing the cause of an unexpected early termination of a science run on June 22.

A real-time procedure was completed on July 15 to update the on-board ephemeris.

The command loads for the coming week include commanding for the first three eclipses of the summer 2004 eclipse season. The first of these eclipses, on July 18, is a penumbra-only eclipse and requires an off-point of the solar arrays to discharge the batteries sufficiently to avoid an over-voltage condition on eclipse exit. Software patches to provide the appropriate power system configuration in the event of spacecraft safing actions during the eclipse will be uplinked and the event monitored in real time on July 18.

The observing schedule for next week is shown below:

------------------------------------------
Radiation Belts                      Jul 18 
LBQS 2358+0216          ACIS-S
Cygnus OB2 North W      ACIS-I
HZ43                    HRC-I/LETG   Jul 19 
HZ43                    HRC-S/LETG
70 Oph                  HRC-S/LETG
F1450-5276              ACIS-S       Jul 20 
Radiation Belts
PSR J1809-1917          ACIS-S       Jul 21 
2QZ J113900.5-0201      ACIS-S
SNR1987A                ACIS-S
Abell 496               ACIS-S       Jul 22 
Radiation Belts                      Jul 23 
SCULPTOR DWARF SPH      ACIS-S       Jul 24 
HH 300                  ACIS-I
RX J105343+5735         ACIS-S
F1410-4555              ACIS-S       Jul 25 
A2151                   ACIS-I

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports